# Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows

### v1

@@ 2.2 Limitations of Existing Optimizations
C2: Joint runtime reconfiguration under concurrency. Beyond per-input routing overheads, runtime reconfiguration amplifies the coordination that serving platforms must consider when managing concurrent requests. Specifically, capitalizing on each request’s routing flexibility can constrain other requests. For instance, consider two requests where R1 can achieve target accuracy with either 7B or 14B models, while R2 needs either 7B or 32B. If the serving engine for the 7B model can only house one more request, independent scheduling might grant R1 use of the 7B model first (its cheapest option), forcing R2 to use the expensive 32B model. Joint scheduling would recognize that R1 can make a slight performance sacrifice by using 14B, allowing R2 to take the scarce 7B slot—improving overall system throughput and latency. However, such joint optimization becomes computationally expensive as it scales exponentially with the number of concurrent requests and configurations.
@@ 3 Design
We present Aragog, an efficient serving system for agentic workflows that dynamically selects workflow configurations and progressively schedules their execution based on real-time system state. Its overarching goal is to optimize the serving capacity of agentic workflows while preserving accuracy. To realize that, the key insight is that the accuracy of a configuration for a given input remains constant, while its runtime cost rapidly fluctuates with system dynamics (i.e., resource availability). Traditional monolithic routers jointly evaluate both accuracy and cost, necessitating frequent and expensive router invocations to account for runtime cost. By recognizing this fundamental asymmetry between accuracy (static and expensive) and cost (dynamic and cheap), we decouple them to make frequent runtime reconfiguration practical: the expensive accuracy evaluation is performed once upfront to identify the viable configuration set, while the lightweight cost-based selection happens frequently at execution time when actual costs are observable. This enables tractable per-input configuration selection (C1) and joint optimization across concurrent requests (C2) at scale.
Specifically, Aragog decouples configuration selection into two components: (1) A configuration predictor3.1) that identifies all accurate configurations for each input before execution, rather than a single configuration. (2) A runtime scheduler3.2) progressively selects from these candidates stage-by-stage at execution time, efficiently coordinating configuration selection across concurrent requests to maximize serving capacity while preserving each request’s flexibility. Figure 5 overviews Aragog’s workflow. 1 Follow the existing systems, Aragog trains configuration predictors offline using representative data. 2 At runtime, incoming queries pass through the configuration predictor to obtain their viable configuration set. 3 Meanwhile, the runtime scheduler monitors the available resources in each model’s serving engine. 4 With this runtime information, the scheduler performs stage-wise scheduling, jointly selecting configurations across all concurrent requests from their respective configuration sets based on slot availability as each request progresses through stages. 5 After each stage, Aragog prunes configurations with different prefixes, progressively narrowing viable options.
Figure 5: Aragog predicts a set of accurate configurations for each input before execution and performs stage-wise scheduling at runtime. At stage 0 example, model C is overloaded, forcing query 2 to use model A. To accommodate this, the scheduler exploits query 1’s flexibility, picking configurations that start with model B.
@@ 3.1 Efficiently Predicting Accurate Configuration Sets
To capitalize on the runtime flexibility of agentic workflows, Aragog’s configuration predictor fundamentally differs from traditional routers–instead of selecting a single best configuration, it must identify the set of accurate configurations for a given input $x$. This configuration set, denoted $\mathcal{C}(x)$, is essential for runtime scheduling flexibility: the prediction must be cheap enough to avoid becoming a bottleneck, while remaining accurate and presenting the maximum number of configurations to maximize runtime flexibility.
Problem formulation. Given an input $x$ and a workflow DAG $G$ with $n$ agents, let $\mathcal{M}=\{m_{1},...,m_{k}\}$ denote the set of candidate models ordered by size (e.g., Qwen-7B through Qwen-32B), where $\mathcal{C}_{\mathrm{all}}=\mathcal{M}^{n}$ represents the full configuration space. We define $c^{*}$ as the most expensive configuration where all agents use the largest model $m_{k}$. Our objective is to find as many configurations as possible whose accuracy matches $c^{*}$ for input $x$, while minimizing overhead.
Decomposing into binary decisions. Existing routing approaches use a standard two-layer architecture: an embedding model (often transformer-based) encodes the input into a high-dimensional representation, followed by a classifier that predicts the configurations to select from $|\mathcal{M}|^{n}$ options. To practically deploy routers, existing systems navigate the tradeoff between router complexity and accuracy. However, identifying the set of viable configurations introduces a fundamentally different problem that opens a larger design space beyond single-router complexity tradeoffs. The design space spans a spectrum of granularities based on how we partition the configuration space. At one extreme, a single router evaluates all $|\mathcal{M}|^{n}$ configurations simultaneously; at the other, $|\mathcal{M}|^{n}$ binary routers each evaluate one configuration independently. Between these extremes lie intermediate approaches, such as $k$ routers each handling $|\mathcal{M}|^{n}/k$ configurations. The fundamental tradeoff is that larger subsets reduce router count but increase per-router complexity, making it difficult to reason about total routing costs.
@@
To systematically explore this tradeoff, we evaluate how the overall routing accuracy and cost accuracy scale with subset size under a fixed parameter budget. We experiment with three workflows, and follow the router architecture in (§5.1). We partition these 16 configurations into different subset sizes: 16 routers handling 1 configuration each (binary), 4 routers handling 4 configurations each, and 1 router handling 16 configurations. As the subset size increases, we scale router parameters linearly to maintain constant total router parameters. Figure 6 shows the results averaged across inputs: Despite allocating the same total parameters across all approaches, the overall accuracy degrades with increasing subset size. Binary routers excel in both precision (fewer inaccurate configurations) and recall (more viable configurations for runtime scheduling). A collection of binary routers achieves 95–97% precision and 47-67% recall, while a single 16-configuration router with the same total parameters achieves only 21–40% precision and 11-21% recall.
This accuracy degradation occurs because routers must learn increasingly complex decision boundaries to distinguish multiple configurations simultaneously, which makes the collection of binary routers the most accurate with the same parameter budget. Therefore, Aragog adopts the finest granularity: decomposing the routing problem into a collection of binary routers $\{f_{c}\}_{c\in\mathcal{C}_{\text{pool}}}$, where each router $f_{c}$ makes a simple binary decision–is configuration $c$ as accurate as $c^{*}$ for input $x$.
@@
Figure 6: Overall routing accuracy drops as each router’s configuration space increases. Binary routers achieve the best precision and recall for a given parameter budget. Our evaluation (§5) confirms that the recall is sufficient to provide flexible runtime scheduling.
With the offline trained binary routers, our goal is for routing to complete during queuing for the first stage of a given request. This requires router inference to be deadline aware–finishing within the request’s estimated queuing delay, as any computation beyond this threshold directly adds latency to that request’s end-to-end serving latency. However, naively running all routers for every possible configuration would itself exceed typical queuing delays. Fortunately, we observe that significant redundancy exists in the configuration space, which we exploit as a knob to dynamically adjust routing inference budget based on estimated queuing delays.
@@
Pruning the configuration space. To reduce the router’s search space, we leverage a key empirical property of model capabilities. We observe that the accuracy of a workflow exhibits near-monotonicity under model upgrades: replacing any agent’s model with a more capable one (i.e., one with more parameters within the same model family) maintains or improves end-to-end accuracy in the vast majority of cases. Figure 7 validates this property empirically across all configurations of a self-refine workflow for the coding task, where green arrows indicate accuracy-preserving upgrades and red arrows indicate rare violations (only 8 out of 288 upgrade cases, with accuracy drops capped at 1.5%). We observe a similar trend consistently across different model families and agentic workflows in our evaluation. The intuition is straightforward: if we run a router on a configuration and it deems that configuration accurate enough, then any configuration that upgrades one or more agents to more capable models are expected to be accurate. Conversely, if a configuration is deemed inaccurate, downgrading any agent will likely remain inaccurate. Built on that, Aragog can effectively prune the whole configuration search space. Crucially, Aragog leverages this empirically observed (though non-guaranteed) property solely to prune the set of configurations; all selected configurations still run through the corresponding binary router before being considered during serving. Thus, violations of this property can only result in reduced serving flexibility by missing viable configurations, as inaccurate configurations will be filtered.
Concretely, the near-monotonic property allows us to avoid exhaustive router evaluation through a two-phase approach: targeted binary search followed by verification. We first construct the search space as a collection of monotonic configuration chains. Each chain starts from the base configuration (all agents using the cheapest model $m_{1}$). We construct these chains through depth-first search (DFS) by individually upgrading each agent’s model while keeping others fixed. Along each chain, we locate the transition boundary between inaccurate and accurate configurations via binary search. The boundary implies that subsequent upgrades are accurate (and all prior downgrades are inaccurate). We add the accurate configurations to our candidate set. This systematic search identifies all boundary configurations without exhaustive router evaluations, dramatically reducing our search from $|\mathcal{M}|^{n}$ total configurations to a much smaller candidate set. Again, before passing these candidates to the scheduler, we run their corresponding binary routers to confirm viability and filter out any false positives.
@@
Efficient router inference Even after pruning, naively evaluating remaining binary routers sequentially is inefficient. As described before, our router architecture follows the standard design used in existing systems: an embedding model (generates a general high-dimensional representation for each input) followed by task-specific classifiers (classify whether specific configurations are accurate). Though the general embedding model can be shared to maximize GPU utilization through batching. The problem is that the sequential execution of specialized classifiers wastes GPU resources–each head’s computational footprint cannot fully utilize GPU resources. To address this, we fuse classification heads into a single batched operation layer by layer through customized GPU kernels. This transforms $k$ sequential memory-bound operations into one compute-bound operation that fully utilizes the GPU compute and reduces kernel launch overheads.
@@
Optimization effectiveness. Figure 8 shows the cumulative impact of our optimizations on average per-input routing latency for two representative workflows. Starting from naive exhaustive evaluation of $|\mathcal{M}|^{n}$ routers (baseline), the monotonic property reduces routing overhead by 30.0–47.6% by configuration pruning. Kernel fusion for classification heads further reduces latency by 33.1–56.3%. Together, these optimizations reduce average per-input routing latency from 90.9–196.2 ms to 42.6–44.9 ms per request.
Putting it all together: deadline-aware configuration search. Aragog adapts its search strategy to the expected budget–the time before the request’s first stage. The scheduler maintains exponential moving averages of per-request queuing times and sets the search deadline accordingly. When queues are deep (high load), longer wait times allow comprehensive search: we perform the full monotonic exploration to identify boundary configurations, then evaluate all remaining candidates with the corresponding binary classification routers. When queues are shallow (low load), we employ early stopping and terminate the search even if more configurations remain unexplored.
@@ 3.2 Runtime Configuration Selection and Scheduling
Given the set of viable configurations from §3.1 for each request and a set of model serving engines (with provisioned resources for each model that Aragog takes as input), Aragog’s goal is to leverage the configuration flexibility and make fine-grained, frequent configuration decisions to best adapt to system dynamics. To realize this, Aragog continuously monitors the available slots for each model’s serving engine. With this fresh view of system state and each request’s viable configurations, Aragog employs stage-wise scheduling that defers configuration selection until necessary, ensuring that selected configurations optimize for the actual resource conditions they encounter to minimize staleness. Moreover, Aragog jointly optimizes across multiple requests to mitigate the contention between concurrent requests, while bounding the coordination overheads (C2).
@@
More specifically, with the following inputs: (i) $\mathcal{Q}=\{r_{1},r_{2},...,r_{n}\}$, the request queue in FIFO order where each $r_{i}$ is at a specific workflow stage; (ii) $\mathcal{C}(r_{i})$, the set of viable configurations for request $r_{i}$; and (iii) $S_{m}(t)$, the available slots for model $m$’s engine at time $t$ (considering current batch occupancy and maximum batch size). The expected output is an assignment of requests to models for the current scheduling round, with the objective of minimizing end-to-end latency and maximizing throughput while preserving the common FIFO ordering. Since each engine’s maximum batch size represents its saturation point, this objective reduces to maximizing utilization across all serving engines–undersaturated engines waste resources and increase queuing delays for pending requests.
@@
Joint optimization across requests and time. Given the mentioned inputs, Aragog must assign each request in the FIFO queue to a configuration from its viable set. The quality of this assignment is critical: it not only leads to resource underutilization in the current round but also future scheduling flexibility, as each configuration selection prunes the configuration set. However, the assignment space is prohibitively large–exponential to the number of requests and the size of their configuration set–making exhaustive search impractical for real-time scheduling.
@@
Our key insight is that FIFO ordering and shrinking configuration sets together narrow the assignment space for efficient exploration. First, FIFO ordering prioritizes earlier requests, and their assignments constrain the available options for later requests due to shared resources. Second, earlier requests in the queue are typically at later workflow stages with already-pruned configuration sets, leaving them with few viable options. Together, these two factors create a narrow search space: earlier requests have limited options, and once assigned, they further restrict later requests. This makes beam search ideal–it avoids both the intractability of exhaustive search and the myopia of greedy assignment by iteratively extending a small set of promising partial assignments in FIFO order.
Figure 9: Beam search balances quality and efficiency, achieving near-optimal assignments comparable to brute-force while maintaining efficiency close to greedy search. Bars list averages across workloads, with error bars for min-max.
@@
Aragog maintains only the top-$B$ (B is defaulted to 4 and evaluate different values in §5.3) most promising partial assignments at each step, where each partial assignment represents a subset of scheduled requests with their configuration assignments. Partial assignments are ranked by two criteria: (1) weighted utilization—filling model slots proportional to each engine’s profiled throughput (e.g., a 32B model slot contributes less than a 7B slot with same amount of serving resource)—and (2) flexibility preservation as a tiebreaker, the percentage of configuration options preserved for request’s future stages. This balances immediate resource contention with preserving a request’s future flexibility.
@@
At each step, the scheduler expands partial assignments in FIFO order by considering all valid configuration assignments for the next unscheduled request. To avoid blocking, we employ look-ahead FIFO: when a request can not be scheduled due to resource constraints, the scheduler skips it and proceeds to the next request in the queue. This is critical for heterogeneous workflows where requests require different models–strict FIFO would block subsequent requests with flexible configurations behind a constrained request waiting for an overloaded model. Our relaxation preserves FIFO fairness because: (1) skipped requests add no additional delay–they remain blocked regardless of whether subsequent requests are scheduled, and (2) Aragog still iterates through requests in FIFO order in future rounds, ensuring skipped requests are never overtaken by later joint requests requiring the same models.
@@
By retaining only $B$ states throughout the search, Aragog makes joint optimization tractable while finding high-quality assignments. The search terminates when all feasible requests are assigned or no valid extensions exist. To evaluate beam search’s performance, we take periodic snapshots of our evaluation workloads during serving. Figure 9 shows that beam search achieves near-optimal assignment quality–measured by resource utilization and flexibility preservation–comparable to exhaustive search while maintaining efficiency close to greedy approaches (order of magnitude faster than brute-force).
@@
Supporting complex workflows. Agentic workflows extend beyond sequential workflows, often containing parallel branches and complex dependencies. Aragog leverages topological sorting to ensure correct execution order while maximizing parallelism. Before execution, Aragog performs a topological sort on the workflow DAG and pre-computes each agent’s depth—the longest path from that agent to any leaf node. During execution, similar to existing systems [parrot, ayo], Aragog follows a two-level priority scheme: (1) inter-request: FIFO ordering prioritizes earlier requests over later ones, and (2) intra-request: agents are prioritized by depth in descending order to schedule critical-path agents first. For example, given two requests $R_{1}$ (arrived first) and $R_{2}$ (arrived second), where $R_{1}$ has ready agents $A$ (depth=4) and $B$ (depth=2), and $R_{2}$ has ready agent $C$ (depth=5), the scheduling order is: $A$ from $R_{1}$ (highest depth in earlier request), then $B$ from $R_{1}$ (same request), and finally $C$ from $R_{2}$ (later request), despite $C$ having the globally highest depth. With this ordering, Aragog’s optimization applies seamlessly to complex workflows.

### v2

@@ 2.3 Challenges
C2: Joint runtime reconfiguration under concurrency. Beyond per-input routing overheads, runtime reconfiguration amplifies the coordination that serving platforms must consider when managing concurrent requests. Specifically, capitalizing on each request’s configuration flexibility can constrain other requests. For instance, consider two requests where R1 can achieve target accuracy with either 7B or 14B models, while R2 needs either 7B or 32B models. If the serving engine for the 7B model can only house one more request in its batch before resorting to queuing, independent scheduling might grant R1 use of the 7B model first (its cheapest option), forcing R2 to use the expensive 32B model. In contrast, joint scheduling would recognize that R1 can make a slight performance sacrifice by using 14B, allowing R2 to take the scarce 7B slot—improving overall system throughput and latency. However, such joint optimization becomes computationally expensive as it scales exponentially with the number of concurrent requests and their configurations.
@@ 3 Design
We present Aragog, an efficient agentic workflow serving system that progressively adapts each request’s configuration to cater to runtime system dynamics, i.e., resource availability. Its overarching goal is to maximize serving capacity (throughput) while preserving accuracy. To do so, our key insight is that the accuracy of a configuration for a given request remains constant, while its runtime cost (achieved throughput, latency, utilization) rapidly fluctuates with system dynamics. By recognizing this fundamental asymmetry – accuracy (static but expensive to compute) vs. cost (dynamic but cheap to assess) – we decouple them to make frequent runtime reconfiguration practical. Specifically, Aragog embeds (1) a configuration predictor3.1) that efficiently identifies all accurate configurations for each input once before execution, rather than just a single configuration as in prior work, to address C1 from §2.3, and (2) a runtime scheduler3.2) that progressively adapts configurations stage by stage (when actual runtime costs are observable) from only this candidate set, efficiently coordinating configuration selection across concurrent requests to maximize serving capacity (addressing C2).
Figure 5: Aragog predicts a set of accurate configurations for each input before execution and performs stage-wise scheduling at runtime. In this stage 0 example, model C is overloaded, limiting query 2 to model A. To jointly optimize both queries, the scheduler exploits query 1’s flexibility, picking configurations that start with model B.
Figure 5 overviews Aragog’s architecture, which runs as a layer between agentic workflow applications and LLM serving engines. Prior to deployment, 1 Aragog trains configuration predictors using representative data for the target workflow, following existing strategies [routellm, mixllm]. Additionally, Aragog profiles the latency and throughput of serving engines under different batch sizes to estimate engine available capacity and configuration cost at runtime. Once deployed, 2 incoming requests are queued in FIFO order and first pass through the configuration predictor to obtain their set of accurate configurations. To align the stage-wise scheduling with runtime system states, the runtime scheduler periodically polls each serving engine’s load 3. When serving engines have available resources, the scheduler initiates a scheduling round 4: using the profiled latency-throughput profiles and current load observations, it jointly selects which requests to dispatch and which models to use for their current stages such that overall serving throughput is maximized. After each scheduling round, Aragog prunes invalid configurations – i.e., those that would have used a different model than the one selected for the current stage – from each scheduled request’s set of candidates 5.
@@ 3.1 Efficiently Predicting Accurate Configuration Sets
To effectively achieve stage-wise configuration adaptation, Aragog first needs to exploit the configuration flexibility of agentic workflows. To do that, Aragog’s configuration predictor fundamentally differs from traditional routers – instead of selecting a single best configuration, it must identify the set of accurate configurations for a given input. This configuration set is essential to Aragog’s operation: the prediction must be cheap enough to avoid becoming a bottleneck, while remaining accurate and presenting the maximum number of configurations to maximize adaptation flexibility.
Problem formulation. We formalize configuration prediction as follows. Given an input $x$ and a workflow DAG $G$ with $n$ agents, let $\mathcal{M}=\{m_{1},...,m_{k}\}$ denote the set of candidate models ordered by size (e.g., Qwen-7B through Qwen-32B), where $\mathcal{C}_{\mathrm{all}}=\mathcal{M}^{n}$ represents the full configuration space. We define $c^{*}$ as the most expensive configuration where all agents use the largest model $m_{k}$, and a configuration is accurate if it achieves equivalent accuracy to $c^{*}$ for a given input. Following standard practice [routellm, mixllm, daao], we employ learned routers to predict accurate configurations, where each router uses a two-layer architecture: an embedding model encodes the input, and is then followed by a classifier. Routers are trained offline on a labeled dataset in which, for each sample, we execute all configurations and label those matching the accuracy of $c^{*}$ positively.
Routing strategy. The key question is how to design routers (and the routing approach) such that accuracy is preserved but overheads are bounded, e.g., with a limit on the number of model parameters used for routing. This leads to a challenging space of options to consider. At one extreme, a single large router evaluates all $|\mathcal{M}|^{n}$ configurations simultaneously; at the other, $|\mathcal{M}|^{n}$ small binary routers each evaluate one configuration independently. Navigating this space is non-trivial: larger routers are more capable, but each must consider more configurations making its task more difficult and the accuracy-compute tradeoff unclear.
To systematically explore this space, we use three representative workflows to evaluate how overall routing accuracy scales with the number of configurations per router under a fixed total router parameter budget. For each workflow, we consider a space of 16 randomly-selected configurations of Qwen model options (7-32B) and evaluate three partitioning schemes: 16 routers each handling 1 configuration (binary routing), 4 routers each handling 4 configurations, and 1 router handling all 16 configurations, with all routers using the same train-test split. For the single-router case, we use the router described in §5.1. As the number of routers increases, we scale down each router’s width proportionally to maintain a constant total parameter budget.
@@
Figure 6 shows the overall routing accuracy results. Despite allocating the same total parameters across all approaches, the overall accuracy degrades with increasing subset size. Binary routers excel in both recall (i.e., finding more viable configurations for runtime scheduling), and even more importantly, precision (i.e., picking fewer inaccurate configurations). Specifically, binary routers achieve 94–96% precision and 71–77% recall, while a single router (handling 16 configurations) achieves only 31–40% precision and 18-21% recall. Building on this, Aragog adopts binary routers to maximize routing accuracy and focuses on enabling efficient inference of those (many) routers via two core techniques, described in turn.
@@
Figure 6: Overall routing accuracy drops as each router’s configuration space increases. Binary routers achieve the best precision and recall for a given parameter budget.
Pruning router inferences. With the trained binary routers, our goal is for routing to complete during queuing for the first stage of a given request, i.e., to avoid added end-to-end serving latency from configuration adaptation. However, naively running all routers can inflate end-to-end request latency by up to 15.4% across our workloads.
Our strategy to address this is rooted in a key empirical observation: the accuracy of a workflow exhibits near-monotonicity under model upgrades, where replacing any agent’s model with a more capable one (i.e., one with more parameters) maintains or improves end-to-end accuracy in the vast majority of cases. Figure 7 validates this property experimentally across all configurations of a self-refine workflow used for code generation, where green arrows indicate accuracy-preserving upgrades and red arrows indicate rare violations (only 8 out of 288 upgrade cases, with accuracy drops capped at 1.5%); note that we observe a similar trend across all other workloads in our evaluation.
@@
Aragog exploits this observation to carefully prune the configuration search space. Intuitively, if we run a router on a configuration and it deems a configuration accurate enough, then any configuration that upgrades one or more agents to more capable models is expected to be accurate. Conversely, if a configuration is deemed inaccurate, downgrading any agent will likely remain inaccurate. Crucially, Aragog leverages this non-guaranteed property solely to prune the set of configurations, i.e., it affects search efficiency, not accuracy; all selected configurations still run through their corresponding routers before being considered during serving with Aragog. Thus, violations of this property can only result in reduced configuration flexibility by missing accurate configurations.
More specifically, Aragog applies this pruning strategy through a binary search. We first structure the search space into monotonic configuration chains, each starting from the base configuration where all agents use $m_{1}$. These chains are built via depth-first search (DFS) by individually upgrading each agent’s model while keeping the others fixed. Then, along each chain, we use binary search to locate the transition boundary between inaccurate and accurate configurations. At each step in the search, we evaluate the selected configuration by running its binary router. Again, after the binary search completes, Aragog verifies all remaining configurations by running their routers and retains only those predicted to be accurate for our final candidate set.
@@
Efficient router inference. While pruning reduces the number of routers to evaluate, executing them efficiently requires further optimization. The core issue is that the specialized classifiers from different routers cannot be batched like the shared embedding model—each classifier is router-specific, resulting in sequential execution. Worse, these classifiers are lightweight, leading to underutilized GPU resources. To address this, Aragog fuses the classifiers layer by layer via custom kernels to maximize the parallelism.
Figure 8 shows the cumulative impact of our optimizations on average per-input routing latency for two representative workflows. Starting from naive exhaustive evaluation of $|\mathcal{M}|^{n}$ routers with embedding batching enabled (baseline), Aragog’s configuration pruning optimization reduces routing overheads by 30.0–47.6%. Kernel fusion for router classifiers further reduces latency by 33.1–56.3%. Together, these optimizations reduce average per-input routing latency from 90.9–196.2 ms to 42.6–44.9 ms per request; we show in §5.3 that the resultant overheads sufficiently ensure that routing minimally affects end-to-end response times.
Overall operation. When deployed online, Aragog adapts configuration search to the time budget – i.e., the configuration search for a request must complete before its first stage can be executed. Specifically, Aragog estimates request queuing delays via exponential moving averages of observed per-stage queue times across all requests. Based on this estimate, Aragog adjusts the embedding model batch size and classifier fusion degree to maximize router inference efficiency while meeting the deadline. If configuration search is incomplete when a request must execute, Aragog uses the verified configurations found so far, or defaults to the most expensive configuration if none exist.
@@ 3.2 Runtime Configuration Selection and Scheduling
Given the set of viable configurations per request (§3.1) and serving engines with provisioned resources to manage all model options, Aragog’s goal is to maximize overall serving capacity. Key to this is ensuring that stage-wise configuration decisions factor in the latest information about resource availability and also jointly optimize across all concurrent requests.
To realize this, Aragog’s scheduler works as follows. It periodically polls each engine’s current load and estimates available capacity relative to offline-profiled saturation points, i.e., the batch size for that model and serving engine beyond which throughput no longer improves. When any engine has available capacity, the scheduler triggers a scheduling round. All requests awaiting scheduling are held in a meta-queue organized in FIFO order (as in existing systems [ayo]), with each request being tagged with both its (remaining) viable configurations and its current workflow stage. With this fresh view of system state and configuration options, each scheduling round involves (1) jointly selecting requests from the queue and assigning them models to use for their current stages, and (2) dispatching those stage-wise inference jobs to the corresponding serving engines which each manage their own queues and execution logic (e.g., chunked prefill decisions). Once dispatched, requests remain in the meta-queue as pending rather than re-entering upon stage completion to preserve FIFO ordering.
@@
Joint optimization across requests and time. Aragog’s runtime scheduler operates using the following inputs for each scheduling round: (i) $\mathcal{Q}=\{r_{1},r_{2},...,r_{n}\}$, the request queue in FIFO order where each $r_{i}$ is at a specific workflow stage; (ii) $\mathcal{C}(r_{i})$, the set of viable configurations for request $r_{i}$; and (iii) $S_{m}(t)$, the available slots for model $m$’s engine at time $t$ (considering current batch occupancy and maximum batch size, as per the offline profiles). In selecting request stages to schedule (and the models to use for each), Aragog strives to maximize overall serving throughput.
@@
This involves not only maximizing current utilization across all serving engines, but also preserving future scheduling flexibility. The reason is that each scheduling round’s configuration assignment has two effects: it determines resource utilization in the current round, and it potentially prunes the set of viable configurations for future stages of the considered requests. Put differently, a model selection for a given stage can render previously viable configuration options invalid (if they conflict at that stage), thereby limiting future adaptation flexibility and (potentially) compromising future resource utilization. Aragog must therefore jointly optimize across requests and time. However, the assignment space is prohibitively large – exponential in the number of requests and the size of their configuration sets – making exhaustive search impractical for real-time scheduling.
@@
To handle this, our key insight is that FIFO ordering and shrinking configuration sets together narrow the assignment space for efficient exploration. First, FIFO ordering prioritizes earlier requests in the queue, and their assignments constrain the available options for later requests due to shared resources. Moreover, earlier requests in the queue are typically at later workflow stages with already-pruned configuration sets, leaving them with few viable options. Together, these two factors create a narrow search space: earlier requests have limited options, and once assigned, they further restrict later requests. This makes beam search ideal – it avoids both the intractability of exhaustive search and the myopia of greedy assignment by iteratively extending a small set of promising partial assignments in FIFO order.
Figure 9: Beam search efficiently navigates the exponential configuration space by iteratively scheduling requests in FIFO order. Given current engine status, at each step it explores configuration options to extend the partial assignment, ranks them by utilization improvement (U) and flexibility preservation (F), and retains only the top-B (B=2 here) for further exploration.
Figure 10: Comparing scheduling algorithms in Aragog: beam search achieves near-optimal resource utilization (current engine utilization) and flexibility preservation (future configuration options) comparable to brute-force, with greedy-level efficiency. Bars show averages with min-max error bars.
@@
Aragog uses beam search to efficiently navigate the exponential configuration space (Figure 9). Aragog maintains only the top-$B$ (B is set to 4 by default, but we evaluate different values in §5.3) most promising partial assignments at each step, where each partial assignment represents a subset of scheduled requests with their configuration assignments. Partial assignments are ranked by two criteria: (1) utilization improvement, i.e., the total forecasted throughput improvement from the assignment, estimated using offline-profiled throughput-load curves for each model engine. This metric accounts for both each engine’s inherent throughput characteristics and its current utilization. For example, the scheduler prefers assignments that distribute load across a lightly-loaded 32B engine and a 7B engine with high marginal throughput gains, rather than assignments that further saturate an already near-capacity 7B engine, and (2) flexibility preservation as a tiebreaker, i.e., the average percentage of configuration options preserved across requests for future stages. This balances immediate resource utilization with flexibility preservation for downstream scheduling.
@@
At each step, the scheduler expands partial assignments in FIFO order by considering all valid configuration assignments for the next unscheduled request in the queue. To avoid blocking, we employ a look-ahead mechanism (that still respects FIFO priorities): when a request cannot be scheduled due to resource constraints, the scheduler skips it and proceeds to the next request in the queue. This is critical for heterogeneous workflows where requests require different models; strict FIFO would block subsequent requests with flexible configurations behind a constrained request waiting for an overloaded model. Our relaxation preserves FIFO fairness because: (1) skipped requests add no additional delay, i.e., they remain blocked regardless of whether subsequent requests are scheduled, and (2) Aragog still iterates through requests in FIFO order in future rounds, ensuring skipped requests are never overtaken by later joint requests requiring the same models.
@@
By retaining only top $B$ partial assignments throughout the search, Aragog makes joint optimization tractable while finding high-quality assignments. The search terminates when all feasible requests are assigned or no available serving resources remain. To evaluate the performance of beam search in Aragog, we take periodic snapshots of our evaluation workloads during serving. Figure 10 shows that beam search achieves near-optimal assignment quality (measured by resource utilization and flexibility preservation) that is comparable to exhaustive search while maintaining efficiency close to greedy approaches, i.e., orders of magnitude faster than brute-force search.
@@
Supporting complex workflows. Thus far, we have assumed sequential workflows. However, agentic workflows extend beyond sequential workflows, often containing parallel branches and complex dependencies. Aragog leverages topological sorting to ensure correct execution order while maximizing parallelism. Before execution, Aragog performs a topological sort [parrot, ayo] on the workflow DAG and pre-computes each agent’s depth, i.e., the longest path from that agent to any leaf node. During execution, Aragog follows a two-level priority scheme: (1) inter-request: FIFO ordering prioritizes earlier requests over later ones, and (2) intra-request: agents are prioritized by depth in descending order to schedule critical-path agents first. For example, given two requests $R_{1}$ (arrived first) and $R_{2}$ (arrived second), where $R_{1}$ has ready agents $A$ (depth=4) and $B$ (depth=3), and $R_{2}$ has ready agent $C$ (depth=5). Agent $A$ from $R_{1}$ (highest depth in earlier request) will be scheduled earlier than agent $C$, despite agent $C$ having the globally highest depth. With this ordering, Aragog’s optimization applies seamlessly to complex workflows.